# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v2

@@
Abstract.
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, through our observations, a globally effective trained model may performance disparities in different clients. This implies that the jointly trained models by clients may lead to unfair outcomes. On the other hand, relevant studies indicate that the transmission of gradients or models in federated learning can also give rise to privacy leakage issues, such as membership inference attacks.
To address the first issue mentioned above, we propose a fairness-aware federated learning algorithm, termed FedFair. Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above. In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility. Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy. Lastly, extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. Code and data is accessible at https://anonymous.4open.science/r/FedFDP-5607.
Federated Learning, Differential Privacy, Fairness
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Federated Learning (FL) (McMahan et al., 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets. After training their local models, each client transmits only the model parameters, not the raw training data, to a central server. The server aggregates these parameters to update the global model, which is then sent back to clients for further training. Thanks to its ability to overcome data silos by jointly improving model performance without the need to upload raw data, FL has gained significant attention in recent years (Zhu et al., 2021)(Lee et al., 2022). In particular, many studies have made great breakthroughs in model performance (Wang et al., 2020; Li et al., 2020b, 2021b) and communication costs (Wang et al., 2019; Ling et al., 2023) in FL.
Figure 1. The test accuracies of 50 clients utilizing a globally trained model in FL.
[…]
@@
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al., 2021c). Current machine learning algorithms that protect fairness, such as  (Dwork et al., 2012; Cotter et al., 2019), have not adequately addressed the issue of performance disparities among clients in federated learning. As shown in Figure 1, the red dots represent the test accuracies of the global model on different clients after completing federated learning with the FedAvg algorithm (McMahan et al., 2017). The global model trained collaboratively by clients exhibits significant unbalanced performance across different datasets held by the various clients. An increasing number of scholars are beginning to pay attention to this performance disparities phenomenon and are starting to work on it  (Li et al., 2020c, 2021b; Zhao and Joshi, 2022).
On the other hand, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data (Zhu et al., 2019) or the membership inference attack (Song et al., 2017). Moreover, Melis et al. (Melis et al., 2022) demonstrated that clients’ training data can be leaked through shared models. In order to enhance data privacy in FL, numerous scholars have begun integrating techniques such as differential privacy (DP) (Dwork et al., 2014), homomorphic encryption (Paillier, 1999), and secure multi-party computation (Yao, 1982) into the federated learning framework. Among these, DP stands out as the premier approach to privacy preservation in federated learning, attributed to its modest computational overhead and robust mathematical underpinnings. This is particularly relevant in defending membership inference attacks, as several studies (Abadi et al., 2016; Carlini et al., 2019; Feldman, 2020) have shown that incorporating appropriate DP during the training phase can effectively prevent inadvertent disclosure of sensitive training information.
Although numerous articles have addressed fairness and differential privacy individually within the context of federated learning, there is a scarcity of research that integrates these two aspects. Achieving a balance between model performance, fairness, and differential privacy protection within federated learning frameworks poses a significant challenge. This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning. By analyzing the relationship between them, we propose a fairness-aware federated learning with differential privacy, effectively guaranteeing model performance when achieving fairness and privacy protection.
Our main contributions are listed as follows:
  1. (1)
    Firstly, a fairness-aware federated algorithm is proposed, denoted as FedFair. It incorporates a novel loss function, designed to optimize the objectives of fairness and model performance simultaneously (Section 4). FedFair algorithm achieves the excellent fairness, as shown by the green dots in Figure1.
  2. (2)
@@
    Building upon FedFair, an algorithm referred to as FedFDP is proposed to achieve differential privacy protection (Section 5). Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1). Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
  3. (3)
    Furthermore, it is demonstrated that an optimal fairness parameter $\lambda$ is found, which, when determined, enables the fastest convergence of FedFDP in balancing model performance and fairness through the convergence analysis (Section 6).
  4. (4)
    And then, a privacy loss analysis is conducted on algorithm FedFDP using the notion of Rényi differential privacy (RDP) to verify its satisfaction of differential privacy requirements (Section 7).
  5. (5)
@@
    Lastly, through a comprehensive empirical evaluation on three public datasets, it is confirmed that FedFair and FedFDP outperform the state-of-the-art solutions in terms of model performance and fairness (Section 8).
2. Related Work
@@
Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning. Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Ullah et al., 2023; Zhou et al., 2023; Fu et al., 2022; Ling et al., 2023; Kerkouche et al., 2021; Fu et al., 2024). In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders. For example, concepts like Good-Intent Fairness (Mohri et al., 2019) and Selection Fairness (Zhou et al., 2021) have been introduced. Further discussions on these concepts can be found in additional literature (Cong et al., 2020; Lyu et al., 2020; Mehrabi et al., 2021; Shi et al., 2021).
@@
At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race. Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al., 2022). There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al., 2021; Padala et al., 2021; Gu et al., 2022, 2020; Pentyala et al., 2022).
@@
The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al. (Li et al., 2020c). This approach ensures fairness by assessing the uniformity of model performance across participating clients. The concept has gained broad acceptance. Building on this, Zhao et al. (Zhao and Joshi, 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy. Concurrently, Li et al. (Li et al., 2021b) concentrated on enhancing the fairness of local models within personalized FL. Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
In comparison to definitions such as group fairness, balanced performance fairness is more meaningful for data providers such as enterprises and institutions, moreover, it can effectively motivate clients participate in the federated system. However, achieving balanced performance fairness typically requires adjusting the direction of model updates. It is a challenge to refine the model for balanced performance fairness while integrating DP techniques, such as clipping and noise addition, into the model’s update process. To the best of our knowledge, we are the first to consider the work of balanced performance fairness in DPFL.
3. Preliminaries
Table 1. Summary of main notations
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
3.1. Federated Learning and Fairness
@@
(1) $\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$
@@
We have defined balanced performance fairness in FL as follows. Specifically, this type of fairness is defined based on the performance variance of the model across clients: the smaller the variance, the fairer the model.
Definition 0.
(Balanced performance fairness (Li et al., 2020c)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
3.2. Differential Privacy
@@
Definition 0.
(Differential Privacy (Dwork et al., 2014)). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
(2) $\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$
@@
Definition 0.
($\mathbf{l_{k}}$-Sensitivity(Dwork et al., 2006)) For a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, we define its $l_{k}$ norm sensitivity (denoted as $\Delta_{k}f$) over all neighboring datasets $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ differing in a single sample as
(3) $\text{sup}_{D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}}||f(D_{i})-f(D_{i}^{\prime})||_{k}\leq\Delta_{k}f.$
DPSGD. In gradient-based methods, DPSGD(Abadi et al., 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  1. (1)
    Firstly, computing the gradient of loss function:
    (4) $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}).$
  2. (2)
    Secondly, clipping the gradient of per-sample to get the sensitivity:
    (5) $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$
    where $C$ represents the clipping bound.
  3. (3)
    Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
    (6) $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$
@@
Output: the final trained model $\mathbf{w}_{T}$
1 Initialize $\mathbf{w}_{0},F(\mathbf{w}_{0})=\text{Initial()}$();
2 SERVER
3       for $t\in[0,1,\cdots,T-1]$  do
4             for $i\in[1,2,\cdots,N]$ parallel  do
6
7            $\mathbf{w}_{t+1}=\sum_{i=1}^{N}p_{i}\mathbf{w}_{t+1}^{i}$;
8             $F(\mathbf{w}_{t+1})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t+1}^{i})$;
9
10      return $\mathbf{w}_{T}$;
11
12LOCAL
14
15      $\mathcal{B}_{i}\leftarrow$ Split $D_{i}$ to batches;
16       for batch $b\in\mathcal{B}_{i}$ do
17             $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i},b)-F(\mathbf{w}_{t})$;
18             $\eta_{i}=\eta\cdot\left(1+\lambda\cdot\Delta_{i}\right)$;
20
22       return $\mathbf{w}_{t+1}^{i},F_{i}(\mathbf{w}_{t+1}^{i})$;
23
Algorithm 1 FedFair
4. Proposed Algorithm: FedFair
@@
(7) $\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
@@
(8) $\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
@@
(9) $\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in section 6.2. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (9), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (9) as follows:
(10) $\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$
@@
(11) $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$
@@
Figure 2. FedFDP framework.
[…]
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Algorithm 1.
  • Server (lines 1-1): The server executes $T$ rounds of communication, each round sending the aggregated model $\mathbf{w}_{t}$ and the loss $F(\mathbf{w}_{t})$ to clients.
  • Local (lines 1-1): For each batch $b$, we compute the adaptive learning rate $\eta_{i}$ in lines 1-1. For line 1, we execute gradient descent by using $\eta_{i}$.
Furthermore, as the training process advances, the model progressively prioritizes fairness, leading to a reduction in $\eta_{i}$. In comparison to training methods with a fixed learning rate, learning rate decay can accelerate the convergence of gradient-based methods (Zinkevich, 2003).
5. Proposed Algorithm: FedFDP
In this section, we propose the FedFDP by integrating differential privacy protection base on FedFair. The flow of FedFDP algorithm as shown in Figure 2. The FedFDP algorithm adds the following two additional processes to FedFair. Firstly, in Section 5.1, we present a fair-clipping strategy that not only achieves sensitivity but also allows adjustments to the level of fairness. Furthermore, in Section 5.2, we expand the use of differential privacy protection to the additional loss values that local clients need to upload, suggesting an adaptive clipping method to maximize utility. The overall FedFDP algorithm is described in Section 5.3.
Input: loss function $F(\mathbf{w})$. Parameters: learning rate $\eta$, noise multiplier for gradient $\sigma$, noise multiplier for loss $\sigma_{l}$, fairness hyperparameter $\lambda$, batch sample ratio $q$, original clipping bound $C$.
Output: the final trained model $\mathbf{w}_{T}$

### v3

@@
Abstract
Federated learning (FL) is a new machine learning paradigm to overcome the challenge of data silos and has garnered significant attention. However, federated learning faces challenges in fairness and data privacy. To address both of the above challenges simultaneously, we first propose a fairness-aware federated learning algorithm, termed FedFair. Then based on FedFair, we introduce differential privacy protection to form the FedFDP algorithm to address the trade-offs among fairness, privacy protection, and model performance. In FedFDP, we designed an fairness-aware gradient clipping technique to identify the relationship between fairness and differential privacy. Through convergence analysis, we determined the optimal fairness adjustment parameters to simultaneously achieve the best model performance and fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping method to minimize privacy budget consumption. Extensive experimental results demonstrate that FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. Codes and datasets will be made public after acceptance.
1 Introduction
@@
Federated Learning (FL) (McMahan et al. 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without exposing their respective datasets. After training their local models, each client only transmits model parameters to the central server, rather than the original training data. The server aggregates these parameters to update the global model, which is then sent back to the clients for further training. Due to its ability to improve model performance through collaboration without the need to upload original data, overcoming data silos, FL has received significant attention in recent years (Zhu, Hong, and Zhou 2021; Lee et al. 2022; Chen et al. 2024). Particularly, many studies have achieved significant breakthroughs in terms of model performance (Wang et al. 2020; Li et al. 2020a, 2021b) and communication costs (Wang et al. 2019; Ling et al. 2024) in FL.
Figure 1: Client drift caused by Non-IID.
@@
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al. 2021a). Current fairness-aware machine learning algorithms, such as (Dwork et al. 2012; Cotter et al. 2019), have not fully addressed the issue of client performance disparities in federated learning. As shown in Figure 1, due to non-independent and identically distributed (Non-IID) data, inconsistent client objectives (yellow squares) lead to significant performance imbalances of the collaboratively trained global model (red square) across different datasets held by various clients. An increasing number of scholars are beginning to focus on this phenomenon of unfair performance disparity (Li et al. 2020b, 2021b; Zhao and Joshi 2022).
On the other hand, recent studies have pointed out that training parameters may leak data privacy, such as recovering original data (Zhu, Liu, and Han 2019) or member inference attacks (Song, Ristenpart, and Shmatikov 2017). Even if only models or gradients are uploaded instead of original data, data leakage can still occur (Melis et al. 2022). To enhance data privacy in FL, differential privacy (DP) (Dwork, Roth et al. 2014) has become the preferred method for protecting privacy in federated learning due to its moderate computational overhead and solid mathematical foundation. Incorporating appropriate DP during the training phase can effectively prevent the accidental leakage of sensitive training information (Abadi et al. 2016; Carlini et al. 2019; Feldman 2020).
Although some works have explored fairness and differential privacy in federated learning separately, no prior research has considered them in a unified framework. For the first time, we propose a fairness-aware federated learning with differential privacy. However, balancing differential privacy protection, fairness, and model performance presents a significant challenge. Specifically, there are two main challenges: First, the clipping and noise addition steps in differential privacy can impact fairness and model utility in federated learning. Second, fair federated learning requires uploading loss information to the central server, which incurs additional privacy budget consumption.
In this paper, to address the first challenge, we analyze the relationship between the fairness loss function and differential privacy in federated learning. We design an fairness-aware gradient clipping strategy to balance fairness and differential privacy. This gradient clipping strategy not only meets the requirements for fairness but also satisfies the sensitivity control needed for differential privacy. Furthermore, through convergence analysis, we demonstrate that an optimal adjustment parameter can be found to achieve the best model performance and fairness. To address the second challenge, we design an adaptive clipping method for loss uploading, significantly reducing the privacy budget consumption in this part.
Our main contributions are as follows:
  • First, we propose a fairness-aware federated algorithm named FedFair. It combines a novel loss function aimed at optimizing the objectives of fairness and model performance simultaneously.
@@
  • Based on FedFair, we propose an algorithm called FedFDP to further equip the system with differential privacy. In particular, we designed a fairness-aware gradient clipping strategy that ensures differential privacy and allows for adjustment of fairness in federated learning. In addition, we propose an adaptive clipping method for the additional loss values uploaded by each client to achieve optimal utility.
  • Further, we conducted a convergence analysis of FedFDP and found the optimal fairness parameter $\lambda$ by minimizing the upper bound of convergence. We analyzed the privacy loss of the FedFDP algorithm using the concept of Rényi Differential Privacy (RDP) to verify its compliance with differential privacy requirements.
@@
  • Finally, through comprehensive empirical evaluations on three public datasets, we confirmed that FedFair and FedFDP are superior to existing solutions in terms of model performance and fairness. What’s more, we investigated the impact of clipping bound and noise multiplier in differential privacy on fairness.
2 Related Work
@@
Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning. Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Fu, Chen, and Han 2022; Zhou et al. 2023; Ling et al. 2024; Fu et al. 2024). In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders. For example, concepts like Good-Intent Fairness (Mohri, Sivek, and Suresh 2019) and Selection Fairness (Zhou, Fang, and Hui 2021) have been introduced. Further discussions on these concepts can be found in additional literature (Cong et al. 2020; Lyu et al. 2020) .
@@
At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race. Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al. 2022). There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al. 2021; Padala, Damle, and Gujar 2021; Gu et al. 2022, 2020; Pentyala et al. 2022).
@@
The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al. (Li et al. 2020b). This approach ensures fairness by assessing the uniformity of model performance across participating clients. The concept has gained broad acceptance. Building on this, Zhao et al. (Zhao and Joshi 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy. Concurrently, Li et al. (Li et al. 2021b) concentrated on enhancing the fairness of local models within personalized FL. Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
3 Preliminaries
3.1 Federated Learning and Fairness
@@
$\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (1)
@@
We have defined balanced performance fairness in FL as follows.
Definition 1
(Balanced performance fairness (Li et al. 2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (2)
3.2 Differential Privacy
@@
Definition 2
(Differential Privacy (Dwork, Roth et al. 2014)). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
$\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$ (3)
@@
DPSGD. In gradient-based methods, DPSGD(Abadi et al. 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  • Firstly, computing the gradient of loss function: $g^{i,j}_{t}=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$.
  • Secondly, clipping the gradient of per-sample to get the sensitivity: $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right)$, where $C$ represents the clipping bound.
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier: $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right)$.
@@
4 FedFair: Equipping FL with Fairness
@@
$\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (4)
@@
$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$ (5)
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 2 of Algorithm 2, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (5), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (5) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (6)
@@
$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (7)
@@
Figure 2: FedFDP framework.
Therefore, our approach can be viewed as establishing an adaptive learning rate that allows the model update to achieve both performance and fairness objectives simultaneously. Based on these principles, we introduce the FedFair algorithm, as depicted in Appendix Algorithm 2.
5 FedFDP: Further Equipping FedFair with Differential Privacy
In this section, we present our algorithm, FedFDP, which further integrates FedFair with differential privacy. The flow of FedFDP algorithm is shown in Figure 2. It adds two additional processes to FedFair:
  • A fair-clipping strategy is added to FedFair. This strategy not only achieves sensitivity but also allows adjustments to the level of fairness.
  • An adaptive clipping method is employed when protecting the loss values that clients share with differential privacy, aiming to maximize utility.
Next, we discuss the details of these two strategies.